# Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS–FTIR

### v1

@@ DISCUSSION
The attenuation coefficients of the substance are determined from its complex refractive index, defined as $\tilde{n}=n+ik$, where $n$ and $k$ are the real and imaginary components, respectively. In the electrostatic approximation [65], the volume attenuation coefficient for small particles depends on the wavenumber and the refractive index of the substance forming the particles [66]:
[…]
Supplementary Materials
This PDF file includes:
Supplementary Text
Figs.S1 to S8
Table S1
@@ Acknowledgements
We sincerely thank Jürgen Smoliner, Alois Lugstein and Masiar Sistani from the Institute of Solid State Electronics at TU Wien for letting us use their FTIR spectrometer. We thank Michael Schneider and Paolo Martini from the Institute of Sensor and Actuator Systems at TU Wien for acquiring the SEM images and helping with finite element method simulations, respectively. Special thanks go to the Invisible–Light Labs team for their technical expertise and supportive encouragement, as well as to our colleagues from the MNS group at TU Wien (Institute of Sensor and Actuator Systems) for their valuable insights and collaboration. We also gratefully acknowledge Georg Pfusterschmied for providing the image in Fig. 1A and Mihnea Surdu from the Extreme Environments Research Laboratory at EPFL in Switzerland for his assistance with the aerosol sampling setup.
Funding: This project received funding from the European Commission under grant agreement HORIZON–EIC–2021–TRANSITIONOPEN–101058011 – NEMILIES, the Austria Wirtschaftsservice (AWS SEED, Deep Tech & Physical Sciences project P2422873), the Austrian Science Fund (FWF) under project I 6086–N, and the Novo Nordisk Foundation under grant number NNF22OC0077964 – MASMONADE.
@@
Conceptualization: JTP, JPL, and SS. Methodology: JTP, JH, EŠ, NL, HB, JPL, and SS. Investigation: JTP, JH, EŠ, NL, AG, and HB. Supervision: AG, JPL, and SS. Writing—original draft: JTP and EŠ. Writing—review & editing: JTP, JPL, NL, JH and SS.
@@
Supplementary Materials for
Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS–FTIR
@@
(Dated: April 14, 2025)
This PDF file includes:
Supplementary Text
Figs. S1 to S8
Table S1
@@ A.   NEMS–FTIR spectra of single NPs
The vertical lines observed in the NEMS–FTIR spectra (visible on Fig. 4C) correspond to characteristic wavenumbers associated with specific nanoplastic particles. Some lines correspond to potential minor oxidation or contamination during sample preparation.
@@ B.   NEMS–FTIR spectra of tea bag leachates
The NEMS–FTIR spectra of tea bag leachates (see Fig. 5C) exhibit vertical lines at the following wavenumbers: 1118, 1173, 1257, 1282, 1371, 1437, 1460, 1553, 1642, 1736, 2860, 2930, 3087, and 3300 $\text{cm}^{-1}$.
Figure S1: Coffee ring effect. SEM image of (A) 1.25 ng of PP nanoparticles, (B) 1.25 ng of PS nanoparticles, (C) a PS–PP–PVC mixture mixed in a mass ratio of 1:1:1, containing 5 ng of each component. Image (C) also shows a close–up view of the coffee ring structure.
@@
Figure S4: Micro– and nanofragments released by tea bags during accelerated aging. (A–C) Magnified views of the membrane sampled with water in which a tea bag had been soaked for 15 days and exposed to environmental stress.
@@
Figure S6: Sampling setups. (A) The PIPEJET nanoDispenser, equipped with two cameras and an XY stage, enables precise positioning of the NEMS chips beneath the dispensing capillary for accurate droplet deposition onto the membrane. (B) Schematic representation of the setup, comprising a nebulizer (Portable Aerosol Generation System, PAGS, Handix Scientific Inc., CO, USA) for aerosolizing the liquid into fine droplets, a diffusion dryer (DDU, 570/L, Topas GmbH, Germany), and a NEMS chip holder (Aerosol Flow Adapter, Invisible–Light Labs GmbH, Austria) containing the NEMS chip. The chip holder was placed in an eight–channel filter sampler (FILT, Brechtel Manufacturing Inc., CA, USA), which includes a built–in pump. The sampling flow rate was set to 0.5 L/min for a duration of 5 minutes for each NEMS chip. Sampling is based on the inertial impaction deposition mechanism.

### v2

@@ METHODS
The attenuation coefficients of the substance are determined from its complex refractive index, defined as $\tilde{n}=n+ik$, where $n$ and $k$ are the real and imaginary components, respectively. For a sparsely distributed particle layer, the decadic linear attenuation coefficient is given by:
$\mu_{\text{10}}(\tilde{\nu})=\dfrac{6\pi\tilde{\nu}}{\ln{(10)}}\operatorname{Im}\left\{\dfrac{\tilde{n}^{2}(\tilde{\nu})-1}{\tilde{n}^{2}(\tilde{\nu})+2}\right\}.$ (7)
A detailed derivation of Eq. (7) can be found in the Supporting Information, section C.
Supporting Information
This file includes:
Lists of characteristic peak positions (in wavenumbers) related to nanoplastic particles and tea bag leachates, derivation of the decadic linear attenuation coefficient for sparse particle film, SEM images of the coffee ring effect, NEMS-FTIR spectra of polypropylene and polyvinyl chloride nanoparticles, SEM images of micro- and nanofragments from tea bag leachates, comparison of non-corrected and corrected ATR-FTIR spectrum of nylon tea bag material, comparison between NEMS-FTIR and ATR-FTIR spectra regarding the spectral artifacts, experimental setup for the accelerated aging of the tea bags, sampling setups, IR spectral features of silicon nitride membranes, dependence of relative photothermal responsivity on the IR beam radius, values for deposited and estimated mass of polystyrene nanoparticles along with the standard deviations
@@ Acknowledgements
The authors sincerely thank Jürgen Smoliner, Alois Lugstein and Masiar Sistani from the Institute of Solid State Electronics at TU Wien for letting us use their FTIR spectrometer. The authors gratefully acknowledge Noelia Barrabés Rabanal and the Research Group for Model Catalysis and Applied Catalysis at TU Wien for kindly providing access to their ATR module and supporting our measurements. The authors thank Michael Schneider and Paolo Martini from the Institute of Sensor and Actuator Systems at TU Wien for acquiring the SEM images and helping with finite element method simulations, respectively. Special thanks go to the Invisible-Light Labs team for their technical expertise and supportive encouragement, as well as to colleagues from the MNS group at TU Wien (Institute of Sensor and Actuator Systems) for their valuable insights and collaboration. The authors also gratefully acknowledge Georg Pfusterschmied for providing the image in Figure 1A. The authors gratefully thank Mihnea Surdu from the Extreme Environments Research Laboratory at EPFL in Switzerland for his assistance with the aerosol sampling setup.
Funding: This project received funding from the European Commission under grant agreement HORIZON–EIC–2021–TRANSITIONOPEN–101058011 – NEMILIES and 101202787 EMILIE Booster, the Austria Wirtschaftsservice (AWS SEED, Deep Tech & Physical Sciences project P2422873), the Austrian Science Fund (FWF) under project I 6086–N, and the Novo Nordisk Foundation under grant number NNF22OC0077964 – MASMONADE.
@@
Conceptualization: JTP, JPL, and SS. Methodology: JTP, JH, EŠ, NL, HB, JPL, and SS. Investigation: JTP, JH, EŠ, NL, AG, and HB. Supervision: AG, JPL, and SS. Writing-original draft: JTP and EŠ. Writing-review & editing: JTP, JPL, NL, and SS.
@@
Supporting Information for
Picogram-Level Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS-FTIR
@@
(Dated: July 19, 2025)
@@ A.   NEMS–FTIR spectra of single NPs
The vertical lines observed in the NEMS–FTIR spectra (visible on Figure 4C) correspond to characteristic peak positions (in wavenumbers) associated with specific nanoplastic particles. Some lines correspond to potential minor oxidation or contamination during sample preparation.
@@ B.   NEMS–FTIR spectra of tea bag leachates
The NEMS–FTIR spectra of tea bag leachates (see Figure 5C) exhibit vertical lines at the following wavenumbers: 575, 990, 1116, 1170, 1210, 1235, 1282, 1371, 1437, 1460, 1553, 1642, 1720, 2860, 2930, 3087, and 3300 $\text{cm}^{-1}$.
C.   Derivation of the decadic linear attenuation coefficient for sparse particle film
In the case of sparsely distributed particles with negligible optical interactions, each particle can be considered an independent absorber. For particles much smaller than the wavelength of incident light $(R<<\lambda)$, the absorption cross section is given by [82]:
[…]
where $\tilde{\nu}=1/\lambda$ is the wavenumber, $\tilde{n}(\tilde{\nu})$ the complex refractive index of the particles, and $n_{m}(\tilde{\nu})$ the real refractive index of the surrounding medium.
The particles sit on top of a suspended dielectric film (SiN membrane), which is optically thin compared to the wavelength $\lambda$. Hence, the refractive index of the medium is approximately that of vacuum $n_{m}(\tilde{\nu})\approx 1$, resulting in:
[…]
The decadic volume attenuation coefficient of a particle is given by:
$\mu_{10,v}(\tilde{\nu})=\dfrac{\sigma_{\text{abs}}(\tilde{\nu})}{\ln(10)},$ (10)
which can be converted to the decadic linear attenuation coefficient through division by the particle volume:
[…]
Figure S1: Coffee ring effect. SEM image of (A) 1.25 ng of PP nanoparticles, (B) 1.25 ng of PS nanoparticles, (C) a PS–PP–PVC mixture mixed in a mass ratio of 1:1:1, containing 5 ng of each component. Image also shows a close–up view of the coffee ring structure.
@@ C.   Derivation of the decadic linear attenuation coefficient for sparse particle film
Figure S4: Micro- and nanofragments released by tea bags during accelerated aging. (A–C) Magnified views of the NEMS membrane resonator sampled with water in which a tea bag had been soaked for 15 days and exposed to environmental stress.
Figure S5: Correction effects and spectral artifacts in ATR–FTIR. (A) Comparison of non-corrected and corrected ATR–FTIR spectrum of nylon tea bag (TB) material. Corrections applied include "Extended ATR correction" (standard mode) and "Baseline correction" using OPUS software. The figure demonstrates how the applied corrections alter the relative intensities of spectral peaks. (B) Comparison of NEMS-FTIR spectra of 100 nL and 500 nL nylon tea bag leachates (TBL) with ATR-FTIR spectra of 500 nL TBL and the nylon TB material in the 2800–1730 cm-1 region. CO2 peaks and diamond phonon bands are visible in the ATR–FTIR spectra, whereas these spectral artifacts are absent in the NEMS–FTIR spectra.
@@
Figure S7: Sampling setups. (A) The PIPEJET nanoDispenser (BioFluidix–Hamilton Freiburg, Germany), equipped with two cameras and an XY stage, enables precise positioning of the NEMS chips beneath the dispensing capillary for accurate droplet deposition onto the membrane. (B) The Drop casting Accessory (Invisible-Light Labs GmbH, Austria) for liquid samples. The accessory allows for precise sample deposition and controlled drying under a mild humidity gradient, ensuring consistent localization of the sample within the perforated membrane area. (C) Schematic representation of the setup, comprising a nebulizer (Portable Aerosol Generation System, PAGS, Handix Scientific Inc., CO, USA) for aerosolizing the liquid into fine droplets, a diffusion dryer (DDU, 570/L, Topas GmbH, Germany), and a NEMS chip holder (Aerosol Flow Adapter, Invisible-Light Labs GmbH, Austria) containing the NEMS chip. The chip holder was placed in an eight-channel filter sampler (FILT, Brechtel Manufacturing Inc., CA, USA), which includes a built–in pump. The sampling flow rate was set to 0.5 L/min for a duration of 5 minutes for each NEMS chip. Sampling is based on the inertial impaction deposition mechanism.